# Search for joint multimessenger signals from potential Galactic PeVatrons with HAWC and IceCube

### v1

@@ Abstract
Galactic PeVatrons are sources that can accelerate cosmic rays to PeV energies. The high-energy cosmic rays are expected to interact with the surrounding ambient material or radiation, resulting in the production of gamma rays and neutrinos. To optimize for the detection of such associated production of gamma rays and neutrinos for a given source morphology and spectrum, a multi-messenger analysis that combines gamma rays and neutrinos is required. In this study, we use the Multi-Mission Maximum Likelihood framework (3ML) with IceCube Maximum Likelihood Analysis software (i3mla) and HAWC Accelerated Likelihood (HAL) to search for a correlation between 22 known gamma-ray sources from the third HAWC gamma-ray catalog and 14 years of IceCube track-like data. No significant neutrino emission from the direction of the HAWC sources was found. We report the best-fit gamma-ray model and 90% CL neutrino flux limit from the 22 sources. From the neutrino flux limit, we conclude that the gamma-ray emission from five of the sources can not be produced purely from hadronic interactions. We report the limit for the fraction of gamma rays produced by hadronic interactions for these five sources.
@@ 1 Introduction
Since the discovery of cosmic rays (CRs) more than 100 years ago (Hess, 1912), but where these high-energy particles come from has not yet been well understood. The highest energy charged particles, protons and nuclei, are thought to be extraterrestrial in origin, and their spectrum can be parametrized using a power law with an index of 2.7 up to PeV (or the “knee” in the spectrum), above which the spectrum steepens (Abbasi et al., 2018). The CRs below the knee are thought to originate from sources within the Milky Way galaxy, hinting at the presence of PeV hadron accelerators within the galaxy known as Galactic PeVatrons (Gabici & Aharonian, 2007). Galactic PeVatrons are expected to accelerate particles up to PeV energies through Fermi acceleration and these accelerated particles are expected to be confined within the Milky Way galaxy (Abbasi et al., 2018).
Supernova remnants (SNRs) are a potential class of Galactic PeVatrons. The particles are accelerated by diffusive shock acceleration in the expanding shock. Only 10% of the energy of SNRs is needed to explain the observed CR energy density in the Galaxy (Blandford & Ostriker, 1978). For the most energetic galactic CRs, it is assumed that only young SNRs can accelerate CRs beyond 100 TeV (Bell et al., 2013). The accelerated CRs can escape the SNR and diffuse into the surrounding interstellar medium. They can interact with the surrounding material like molecular clouds and produce charged and neutral pions. Therefore, young SNRs located near star clusters or molecular clouds are thought to be bright in gamma rays due to hadronic interactions. High-energy gamma rays and neutrinos can be efficiently produced by the decay of charged and neutral pions. Other potential source classes of Galactic PeVatrons include young massive star clusters, YMCs (Bykov et al., 2020) and pulsar wind nebulae, PWNe (Di Palma et al., 2017).
@@
Several potential Galactic PeVatrons have been investigated by the community. The HAWC collaboration reported that the TeV gamma-ray spectrum of HAWC J1825-134 extends up to 200 TeV without any cut-off, much higher than the Klein-Nishina suppression energy(Albert et al., 2021). This may indicate hadronic acceleration. The HAWC collaboration also reported HAWC J2227+610 as a potential PeVatron candidate from TeV gamma-ray observations and concluded the proton cutoff energy is at least 800 TeV (Albert et al., 2020a). The HAWC collaboration studied the ultra-high-energy spectrum off MGRO J1908+06 and their model allows a hadronic component at the highest energy (Albert et al., 2022). The HESS Collaboration reported evidence of a PeVatron at the galactic center based on gamma-ray observations up to 40 TeV (Abramowski et al., 2016).
The IceCube Neutrino Observatory has previously searched for neutrino emission from galactic sources, both individual sources and classes of sources. Abbasi et al. (2023a) searched for neutrino emission from LHAASO’s first catalog of sources with energy above 100 TeV (Cao et al., 2021) using a fixed spatial extension and a power law spectrum. They found no significant emission, neither from single sources nor from stacked catalogs of SNRs and PWNe. Recently, IceCube has detected neutrinos from the galactic plane (Abbasi et al., 2023b). The observation was made using a neutrino detection channel, cascades, which has relatively large angular uncertainty. Therefore, IceCube could not distinguish between diffuse emission in the galactic plane and a collection of individual sources. The IceCube collaboration also searched the entire galactic plane and regions near TeV gamma-ray sources using an extended source hypothesis and a power law spectrum, finding no significant neutrino emission (Abbasi et al., 2023c). However, these searches have not directly incorporated gamma-ray data. The simultaneous use of both data sets incorporates more physically motivated spectra and provides a direct constraint on the hadronic ratio, i.e. the fraction of gamma rays originating from hadronic interactions.
@@ 2.1 HAWC Gamma-ray Observatory
In this study, we use 2141 days of HAWC data using the Pass 5 reconstruction (Yun-Carcamo et al., 2023). The event-related energy is estimated using the “ground parameter” estimator described in Abeysekara et al. (2019a). This energy estimator uses the charge at a fixed optimal distance from the shower axis to estimate the energy. The dataset covers the period from June 15, 2015 to October 21, 2021 with a livetime of more than 90%.
@@ 3.1 Source selection
HAWC detected 65 sources with a significance above in the third HAWC catalog (3HWC) (Albert et al., 2020b) using 1523 days of data. The 3HWC provides the location and the gamma-ray flux at a pivot energy of 7 TeV for each source in the catalog. The pivot energy is the energy of normalization of the spectral function, which is specified in terms of . If the gamma-ray emission from these sources comes from hadronic interactions, the neutrino flux will be described by the same spectral function with half the pivot energy (3.5 TeV), according to Equation 2. We calculate the IceCube 90% confidence level (CL) sensitivity for a point source with an unbroken power law spectrum. The sensitivity is defined as the median expected 90% CL upper limit. The sensitivity can be determined using background trials and signal trials. Background trials are done by performing the analysis using simulated data obtained by scrambling real data in right ascension (RA). Scrambling data as background does not suffer from systematic uncertainties and therefore is a robust way to generate background trials. Signal trials are generated by scrambling data and adding signal using simulation data. The calculation of sensitivity is based on a maximum likelihood method as used in the IceCube 10-year search for neutrino sources(Aartsen et al., 2020a). In this work, we use a new likelihood analysis software component called i3mla (Abbasi et al., 2021) with the multi-mission maximum likelihood framework (3ML) (Vianello et al., 2017). Details of i3mla are described in section 3.4. The likelihood consists of a spatial term and an energy term, as explained in more detail in Appendix A.
We select sources whose predicted neutrino flux from hadronic interactions is larger than the sensitivity of the IceCube point source search. The hard spectrum is an optimistic choice made to avoid missing any potential sources in our source selection step. The final sensitivity of the joint search will change depending on the best-fit gamma-ray spectrum and spatial extension. The sensitivity of IceCube and the predicted neutrino flux from 3HWC sources are shown in Figure 1. We exclude two extra-galactic sources, Markarian 421 and Markarian 501, as this study focuses on galactic sources. We also exclude four sources classified as secondary sources in 3HWC, as these are likely to be statistical fluctuations near bright sources. Our final source list comprises 22 sources from 3HWC. Many of the sources are still unidentified but some have potential associations with pulsar wind nebulae, supernova remnants, and star-forming regions.
@@ 3.2 Model selection and fit using gamma-ray data
To obtain a spatial and spectral model for the joint analysis of the sources, we first fit each gamma-ray source with an updated HAWC data set. We use 2141 days of HAWC ground parameter energy estimator maps with the Pass 5 reconstruction for our model selection procedure. The dataset contains more data and a better energy estimator compared to the dataset used in 3HWC. The new reconstruction provides a better angular reconstruction and gamma/hadron separation. We use the 3ML framework (Vianello et al., 2017) with the HAWC Accelerated Likelihood (HAL) plugin (Abeysekara et al., 2021) for our model selection and model fitting. HAL is a Python-based, maximum likelihood software module designed to perform HAWC analysis using a binned likelihood. In conjunction with 3ML, it can perform both stand-alone HAWC analyses and joint analyses with other experiments. We select the model that best describes the gamma-ray emission based on the log-likelihood (LLH). We use the position in 3HWC as a starting point for our initial fit. First, we use a simple power law and a point source as spectral and spatial hypotheses. The simple power law is parameterized as , where is the flux normalization and is the spectral index. The free parameters in this model are the Right Ascension (RA), the Declination (Dec), the spectral index, and the gamma-ray flux. Second, we test the extended source hypothesis, where we fix the RA and Dec of the source to the best-fit values and fit for a Gaussian extension along with the spectral index and flux. We accept the extended source hypothesis if twice the difference between the log-likelihood () of the two models (point source hypothesis and extended source hypothesis) is greater than 16, corresponding to according to Wilks’ theorem (Wilks, 1938). Finally, we test the spectrum curvature hypothesis, where we fixed the RA, Dec, and extension (if the source is extended) and we use a log-parabola spectrum to fit the data again. The log-parabola spectrum is parameterized by
@@ 3.3 IceCube likelihood with i3mla
In this analysis, we use spatial and energy likelihood terms to calculate the likelihood of the model. The signal hypothesis assumes some neutrino emission from the source of interest. The background hypothesis assumes all non-neutrino events and neutrino events originate from atmospheric muons, atmospheric neutrinos, and astrophysical neutrinos that are not associated with the source. We model the IceCube point spread function (PSF) as a 2D Gaussian and the angular error of the 2D Gaussian is estimated during the reconstruction. The signal spatial likelihood is the PSF convolved with the morphology of the source and the signal energy likelihood is calculated using the IRFs and the model spectrum. The background spatial likelihood is estimated from the data by binning the data in reconstructed declination and creating a spline using the resulting histogram. The background spatial likelihood is a function of declination only. The background energy likelihood is estimated from the data by weighting the simulation using the spectral hypothesis and binning it in reconstructed declination and energy. The method is conceptually identical to the previous IceCube 10-year time-integrated point source search in Aartsen et al. (2020a) and the extended source search in Abbasi et al. (2023a, c). A detailed description of the likelihood can be found in Appendix A.
@@ 3.4 Joint fit with Gamma rays and Neutrinos
To calculate the p-value, we perform a large number of background trials by scrambling the IceCube data in RA and building the background test statistic (TS) distribution (Aartsen et al., 2020b). This is done to avoid any dependence on the simulation when testing for a neutrino source. The calculation of the IceCube component of the likelihood is fast. However, using raw HAWC binned maps for computing the HAWC component of the likelihood is prohibitively computationally expensive for such a large number of trials. Therefore, the HAWC data are fixed during the background scrambling process so that we can pre-compute the HAWC component of the likelihood. We extract the HAWC likelihood around the best-fit spectral parameters and save it into a table. Since the location and extension are fixed during the fit, we do not extract the likelihood around the spatial parameters. For spectral parameters, we evaluate the HAWC likelihood within of the best-fit spectral shape parameters while maximizing it over the flux normalization.
Since the gamma-ray and neutrino fluxes are floated independently, maximizing the likelihood over the gamma-ray flux normalization leads to a maximum likelihood for HAWC at the spectral shape parameters in the joint fit. For a log-parabola spectrum, there are two free parameters for the spectral shape ( and ), and we use a 5050 grid around the best-fit parameters for the likelihood evaluation. For a power-law spectrum, there is only one free parameter for the spectral shape () and we evaluate 300 points around the best-fit parameters. We have developed a new custom plugin for 3ML to read the likelihood table and construct a spline. The 3ML plugin returns the spline value for the likelihood evaluation. This method speeds up the likelihood calculation because the program does not have to calculate the HAWC likelihood for each pixel and each bin during the fitting. Figure 2 shows an example of the likelihood.
The background TS distribution is calculated by generating 50,000 background-only trials. For each background trial, the IceCube data are scrambled in RA while keeping the HAWC component of the likelihood unchanged, and the joint analysis is performed. To determine the sensitivity to the neutrino flux of the joint search, we inject different simulated values of the neutrino flux at the location of the best-fit gamma-ray source with its best-fit extension and spectral shape. The 90% CL sensitivity is the neutrino flux required for the TS to exceed the median of the background-only TS distribution 90% of the time. The discovery potential is the neutrino flux required for the TS to exceed the fluctuation of the background-only TS distribution 50% of the time. Figure 3 shows the ratio between the sensitivity and the predicted neutrino flux assuming pp interactions, as well as the ratio between the discovery potential and the predicted neutrino flux assuming pp interactions.

### v2

# Search for joint multimessenger signals from potential galactic cosmic-ray accelerators with HAWC and IceCube

@@ Abstract
The origin of high-energy galactic cosmic rays is yet to be understood, but some galactic cosmic-ray accelerators can accelerate cosmic rays up to PeV energies. The high-energy cosmic rays are expected to interact with the surrounding material or radiation, resulting in the production of gamma rays and neutrinos. To optimize for the detection of such associated production of gamma rays and neutrinos for a given source morphology and spectrum, a multi-messenger analysis that combines gamma rays and neutrinos is required. In this study, we use the Multi-Mission Maximum Likelihood framework (3ML) with IceCube Maximum Likelihood Analysis software (i3mla) and HAWC Accelerated Likelihood (HAL) to search for a correlation between 22 known gamma-ray sources from the third HAWC gamma-ray catalog and 14 years of IceCube track-like data. No significant neutrino emission from the direction of the HAWC sources was found. We report the best-fit gamma-ray model and 90% CL neutrino flux limit from the 22 sources. From the neutrino flux limit, we conclude that for five of the sources, the gamma-ray emission observed by HAWC cannot be produced purely from hadronic interactions. We report the limit for the fraction of gamma rays produced by hadronic interactions for these five sources.
@@ 1 Introduction
Cosmic rays (CRs) were discovered more than 100 years ago (Hess, 1912), but where these high-energy particles come from has not yet been well understood. The highest energy charged particles, protons and nuclei, are thought to be extraterrestrial in origin, and their spectrum can be parametrized using a power law with an index of 2.7 up to PeV (or the “knee” in the spectrum), above which the spectrum steepens (Nagano et al., 1984; Abbasi et al., 2018; Cao et al., 2024a; Amenomori et al., 2008; Aartsen et al., 2020a; Glasmacher et al., 1999). The CRs below the knee are thought to originate from sources within the Milky Way galaxy, hinting at the presence of PeV hadron accelerators within the galaxy known as Galactic PeVatrons (Gabici & Aharonian, 2007). Galactic PeVatrons are expected to accelerate particles up to PeV energies through Fermi acceleration and these accelerated particles are expected to be confined within the Milky Way galaxy (de Oña Wilhelmi et al., 2024). The AS gamma collaboration reported the detection of diffuse gamma rays from the galactic disk with energy between 100 TeV and 1 PeV, hinting at the presence of PeVatrons within our galaxy (Amenomori et al., 2021).
Supernova remnants (SNRs) are a potential class of Galactic PeVatrons. The particles are accelerated by diffusive shock acceleration in the expanding shock. Only 10% of the energy of SNRs is needed to account for the observed CR energy density in the Galaxy (Blandford & Ostriker, 1978). For the most energetic galactic CRs, it is assumed that only young SNRs can accelerate CRs beyond 100 TeV (Bell et al., 2013). The accelerated CRs can escape the SNR and diffuse into the surrounding interstellar medium. They can interact with the surrounding material like molecular clouds and produce charged and neutral pions. Therefore, young SNRs located near star clusters or molecular clouds are thought to be bright in gamma rays due to hadronic interactions. High-energy gamma rays and neutrinos can be efficiently produced by the decay of charged and neutral pions. Other potential classes of Galactic PeVatrons include young massive star clusters, YMCs (Bykov et al., 2020) and pulsar wind nebulae, PWNe (Di Palma et al., 2017).
@@
Several potential Galactic PeVatrons have been investigated by the community. The HAWC collaboration reported that the TeV gamma-ray spectrum of HAWC J1825-134 extends up to 200 TeV without any cut-off, much higher than the Klein-Nishina suppression energy (Albert et al., 2021). This may indicate hadronic acceleration. The HAWC collaboration also reported HAWC J2227+610 as a potential PeVatron candidate from TeV gamma-ray observations and concluded the proton cutoff energy is at least 800 TeV (Albert et al., 2020a). The HAWC collaboration studied the ultra-high-energy spectrum of MGRO J1908+06 and their model allows a hadronic component at the highest energy (Albert et al., 2022). The HESS Collaboration reported evidence of a PeVatron at the galactic center based on gamma-ray observations up to 40 TeV (Abramowski et al., 2016). The LHAASO collaboration reported an ultrahigh-energy gamma ray bubble up to a few PeV from the direction of the Cygnus X region, and suggested that a super PeVatron with maximum proton energy of 10 PeV could power the source (LHAASO Collaboration, 2024). The LHAASO collaboration detected 12 gamma-ray sources with emissions above 100 TeV (Cao et al., 2021) and recently reported their first catalog 1LHAASO (Cao et al., 2024b). The 1LHAASO contains 90 gamma-ray sources, and 43 sources have emission above 100 TeV, making them possible Galactic PeVatrons.
Many efforts have been made to search for PeVatrons using neutrino data. Huang & Li (2022) set a constraint on neutrino emission from LHAASO’s first catalog of sources with energy above 100 TeV (Cao et al., 2021) using previous IceCube neutrino upper limits. They used measurements from other gamma-ray instruments and constrained the neutrino flux using the measured gamma-ray spectrum. Another search by Huang & Li (2022) used 10 years of public IceCube track data to search for neutrino emission from the same LHAASO catalog (Cao et al., 2021) with fixed morphology and different spectral models. The IceCube Neutrino Observatory has previously searched for neutrino emission from galactic sources, both individual sources and classes of sources. Abbasi et al. (2023a) searched for neutrino emission from LHAASO’s first catalog of sources with energy above 100 TeV (Cao et al., 2021) using a fixed spatial extension and a power law spectrum and 12 years of IceCube track data. They found no significant emission, neither from single sources nor from stacked catalogs of SNRs and PWNe. Recently, IceCube has detected neutrinos from the galactic plane (Abbasi et al., 2023b). The observation was made using a neutrino detection channel, cascades, which has relatively large angular uncertainty. Therefore, IceCube could not distinguish between diffuse emission in the galactic plane and a collection of individual sources. The IceCube collaboration also searched the entire galactic plane and regions near TeV gamma-ray sources using an extended source hypothesis and a power law spectrum, finding no significant neutrino emission (Abbasi et al., 2023c). However, these searches have not directly incorporated gamma-ray data. The simultaneous use of both data sets incorporates more physically motivated spectra and more accurate morphology. In this work, for the first time, we simultaneously use both gamma-ray and neutrino data to improve the sensitivity of the search and provide a direct constraint on the hadronic ratio, i.e., the fraction of gamma rays originating from hadronic interactions. The improvement in sensitivity of the search comes from the additional 2 years for data taken with full detector compared to Abbasi et al. (2023a) and the modeling of the neutrino emission under the assumption that the gamma rays and neutrinos originated from the same population of hadrons.
@@ 2.1 HAWC Gamma-ray Observatory
In this study, we use 2141 days of HAWC data using the Pass 5 reconstruction (Albert et al., 2024). The event-related energy is estimated using the “ground parameter” estimator described in Abeysekara et al. (2019a). This energy estimator uses the charge at a fixed optimal distance from the shower axis to estimate the energy. The dataset covers the period from June 15, 2015 to October 21, 2021 with a livetime of more than 90%. The dataset contains 29% more data with an improved reconstruction compared to the dataset used in the third HAWC catalog (3HWC) (Albert et al., 2020b).
@@ 3.1 Source selection
The third HAWC catalog (3HWC) includes 65 sources with a significance above 5 based on 1523 days of data (Albert et al., 2020b). The 3HWC provides the location and the gamma-ray flux at a pivot energy of 7 TeV for each source in the catalog. The pivot energy is the energy of normalization of the spectral function, which is specified in terms of . If the gamma-ray emission from these sources comes from hadronic interactions, the neutrino flux will be described by the same spectral function with half the pivot energy (3.5 TeV), according to Equation 2. We calculate the IceCube 90% confidence level (CL) sensitivity for a point source with an unbroken power law spectrum. The sensitivity is defined as the median expected 90% CL upper limit. The sensitivity can be determined using background trials and signal trials. Background trials are done by performing the analysis using simulated data obtained by scrambling real data in right ascension (RA). Scrambling data as background does not suffer from systematic uncertainties and therefore is a robust way to generate background trials. Signal trials are generated by scrambling data and adding signal using simulation data. The calculation of sensitivity is based on a maximum likelihood method as used in the IceCube 10-year search for neutrino sources(Aartsen et al., 2020b). In this work, we use a new likelihood analysis software component called i3mla (Abbasi et al., 2021) with the multi-mission maximum likelihood framework (3ML) (Vianello et al., 2017). Details of i3mla are described in section 3.4. The likelihood consists of a spatial term and an energy term, as explained in more detail in Appendix A.
We select sources whose predicted neutrino flux from hadronic interactions is larger than the sensitivity of the IceCube point source search. The source selection step is motivated by the trade-off between searching as many possible sources as possible to avoid missing potential sources and avoiding suffering sensitivity loss from the trial factor. The hard spectrum is an optimistic choice made to avoid missing any potential sources in our source selection step since it is close to the hardest spectrum expected. The sensitivity of the joint search will depend on the best-fit gamma-ray spectrum and spatial extension and is different from the point source sensitivity here. The sensitivity of IceCube and the predicted neutrino flux from 3HWC sources are shown in Figure 1. We exclude two extra-galactic sources, Markarian 421 and Markarian 501, as this study focuses on galactic sources. We also exclude four sources classified as secondary sources in 3HWC, as these are likely to be statistical fluctuations near bright sources. Our final source list comprises 22 sources from 3HWC. Many of the sources are still unidentified but some have potential associations with pulsar wind nebulae, supernova remnants, and star-forming regions.
@@ 3.2 Model selection and fit using gamma-ray data
To obtain a spatial and spectral model for the joint analysis of the sources, we first fit each gamma-ray source with an updated HAWC data set. We use 2141 days of HAWC ground parameter energy estimator maps with the Pass 5 reconstruction for our model selection procedure. The dataset contains 29% more data with an improved reconstruction compared to the dataset used in 3HWC. The new reconstruction provides a better angular reconstruction and gamma/hadron separation. We use the 3ML framework (Vianello et al., 2017) with the HAWC Accelerated Likelihood (HAL) plugin (Abeysekara et al., 2021) for our model selection and model fitting. HAL is a Python-based, maximum likelihood software module designed to perform HAWC analysis using a binned likelihood. In conjunction with 3ML, it can perform both stand-alone HAWC analyses and joint analyses with other experiments. We select the model that best describes the gamma-ray emission based on the log-likelihood (LLH). We use the position in 3HWC as a starting point for our initial fit. First, we use a simple power law and a point source as spectral and spatial hypotheses. The simple power law is parameterized as , where is the flux normalization and is the spectral index. The free parameters in this model are the Right Ascension (RA), the Declination (Dec), the spectral index, and the gamma-ray flux. Second, we test the extended source hypothesis, where we fix the RA and Dec of the source to the best-fit values and fit for a Gaussian extension along with the spectral index and flux. We accept the extended source hypothesis if twice the difference between the log-likelihood () of the two models (point source hypothesis and extended source hypothesis) is greater than 16, corresponding to according to Wilks’ theorem (Wilks, 1938). Finally, we test the spectrum curvature hypothesis, where we fixed the RA, Dec, and extension (if the source is extended) and we use a log-parabola spectrum to fit the data again. The log-parabola spectrum is parameterized by
@@ 3.3 IceCube likelihood with i3mla
In this analysis, we use spatial and energy likelihood terms to calculate the likelihood of the model. The signal hypothesis assumes some neutrino emission from the source of interest. The background hypothesis assumes all non-neutrino events and neutrino events originate from atmospheric muons, atmospheric neutrinos, and astrophysical neutrinos that are not associated with the source. We model the IceCube point spread function (PSF) as a 2D Gaussian and the angular error of the 2D Gaussian is estimated during the reconstruction. The signal spatial likelihood is the PSF convolved with the morphology of the source and the signal energy likelihood is calculated using the IRFs and the model spectrum. The background spatial likelihood is estimated from the data by binning the data in reconstructed declination and creating a spline using the resulting histogram. The background spatial likelihood is a function of declination only. The background energy likelihood is estimated from the data by weighting the simulation using the spectral hypothesis and binning it in reconstructed declination and energy. The method is conceptually identical to the previous IceCube 10-year time-integrated point source search in Aartsen et al. (2020b) and the extended source search in Abbasi et al. (2023a, c). A detailed description of the likelihood can be found in Appendix A.
@@ 3.4 Joint fit with Gamma rays and Neutrinos
To calculate the p-value, we perform a large number of background trials by scrambling the IceCube data in RA and building the background test statistic (TS) distribution (Aartsen et al., 2020c). This is done to avoid any dependence on the simulation when testing for a neutrino source. The calculation of the IceCube component of the likelihood is fast. However, using raw HAWC binned maps for computing the HAWC component of the likelihood is prohibitively computationally expensive for such a large number of trials. Therefore, the HAWC data are fixed during the background scrambling process so that we can pre-compute the HAWC component of the likelihood. We extract the HAWC likelihood around the best-fit spectral parameters and save it into a table. Since the location and extension are fixed during the fit, we do not extract the likelihood around the spatial parameters. For spectral parameters, we evaluate the HAWC likelihood within of the best-fit spectral shape parameters while maximizing it over the flux normalization.
Since the gamma-ray and neutrino fluxes are floated independently, maximizing the likelihood over the gamma-ray flux normalization leads to a maximum likelihood for HAWC at the spectral shape parameters in the joint fit. For a log-parabola spectrum, there are two free parameters for the spectral shape ( and ), and we use a 5050 grid around the best-fit parameters for the likelihood evaluation. For a power-law spectrum, there is only one free parameter for the spectral shape () and we evaluate 300 points around the best-fit parameters. We have developed a new custom plugin for 3ML to read the likelihood table and construct a spline. The 3ML plugin then returns the spline values for the likelihood evaluations. This method speeds up the likelihood calculation because the program does not have to calculate the HAWC likelihood for each pixel and each bin during the fitting. Figure 2 shows an example of the likelihood.
The background TS distribution is calculated by generating 50,000 background-only trials. For each background trial, the IceCube data are scrambled in RA while keeping the HAWC component of the likelihood unchanged, and the joint analysis is performed. To determine the sensitivity of the joint search for a source with the best-fit parameters, we inject different simulated values of the neutrino flux at the location of the best-fit gamma-ray source with its best-fit extension and spectral shape. The 90% CL sensitivity is the neutrino flux required for the TS to exceed the median of the background-only TS distribution 90% of the time. The discovery potential is the neutrino flux required for the TS to exceed the fluctuation of the background-only TS distribution 50% of the time. Figure 3 shows the ratio between the sensitivity and the predicted neutrino flux assuming pp interactions, as well as the ratio between the discovery potential and the predicted neutrino flux assuming pp interactions.